A sewing workshop with an annual production volume of 2 thousand units bears the cost of renting premises in the amount of UAH 25 thousand. and in the form of interest payments of UAH 11 thousand per year. In addition, the cost of labor for the release of a unit of the product is 20 UAH, and the cost of raw materials and materials 12 UAH. Calculate total, fixed and variable costs of production, as well as average total costs. What is the difference between total costs and average costs?


1
Expert's answer
2021-02-24T15:03:47-0500

"Q=2000"


"FC=25000+11000=36000"

"VC=AVC\times Q"

"AVC=20+12=32"

"VC=32\times2000=64000"

"TC=FC+VC"

"TC=36000+64000=100000"

"ATC=\frac {TC}{Q}"

"ATC=50"

The total costs cover all costs of the enterprise, and the averages provide information on the costs per unit of output.